Some data from a lab class's free-fall measurements are shown below using Capstone's statistics visualization display.

12 Max: 11.7360 10 Mean: 9.9146 Std. Dev..1.1313 Min.: 8.1640 0 4. 12 Group

Use the graph to calculate

The fraction of data points within 1 standard deviation:

The percent precision of this data set:
Answer #1

Note that there are a total of 13 observations taken.

Mean of these observations = 9.9146 m/s²

And, standard deviation = 1.1313 m/s²

So, one standard deviation interval from mean would be:

I = (9.9146-1.1313, 9.9146+1.1313)

= (8.7833, 11.0459)

So, from graph we can see that there are 8 observations within this interval.

Thus, fraction of data points within one standard deviation = 8/13 = 0.6154.

---------------------------------------------------------------

And the percentage precision is -

Standard Deviation Mean 1.1313 x 100 11.41% 00116 x 100 9.9146